OpenStudy (anonymous):

The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

OpenStudy (anonymous):

Do you know the formula for the equation of a straight line given the gradient and a point on said line?

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

Okay so given a point (x1, y1) and a gradient m. The equation is: (y-y1)=m(x-x1). Our gradient is -4 and our point is (-8,2). Putting these values in we have: y-2=4(x+8) Rearranging this we get: y=4x+34

OpenStudy (anonymous):

oh okay that makes sense. Thank you so much for explaining it!!

OpenStudy (anonymous):

No problem, happy to help!